# Data-Efficient Quantum Noise Modeling via Machine Learning

### v1

@@ III.3 Random circuits
We further benchmark the method through random circuits on ibmq_kolkata. Random quantum circuits were generated using Qiskit’s random_circuit function with systematically varied parameters: qubit count $n\in\{4,...,10\}$ and depth $d=n+3$. The function constructs circuits by randomly sampling single-qubit gates from SU(2) and placing two-qubit gates between random qubit pairs within the depth constraint, with terminal measurements on all qubits. The results show a substantial mean improvement of 43.60% in model fidelity for training and 45.90% for prediction (see Fig. 7(b) in Appendix A). We perform three independent optimization processes and achieve a mean Hellinger distance reduction of $42.9\%\pm 1.5\%$ for the training set and $46.8\%\pm 2.1\%$ for the prediction set, with a maximum reduction of 63% observed for a circuit in the prediction set. This demonstrates that our method’s high performance is a statistically significant result, making it a reliable tool for practical hardware characterization.
@@ IV Discussion and Conclusion
This work develops a novel parameterized noise model with circuit-size-independent complexity. Leveraging machine learning, we repurpose data from a few routine circuit executions to construct high-fidelity, hardware-specific noise models for superconducting quantum processors. By systematically searching complex, multi-parameter error spaces, our approach consistently produces models that are significantly more predictive than default ones obtained from hardware vendors. We demonstrate a reduction of up to 65% in the Hellinger distance, a metric that quantifies the divergence between predicted and experimental quantum state distributions, confirming that our ML-optimized models produce output distributions that more accurately represent the behavior of algorithms executed on real quantum hardware. This significant improvement in model fidelity is essential for the reliable simulation of quantum circuits on noisy quantum devices, enabling a more accurate noise-aware quantum circuit compilation by providing a precise foundation for optimizing qubit selection, SWAP gate insertion, and other compilation passes.
@@
A key advantage of our approach is its efficiency and practicality. It repurposes existing benchmarking or application-specific execution data, avoiding the need for extensive characterization protocols such as quantum process tomography. This reusability reduces experimental overhead, enabling practitioners to leverage prior executions to improve algorithm testing and error mitigation protocols. Moreover, this allows for the continuous refinement of noise models, ensuring scalability for larger quantum circuits and processors. Comprehensive benchmarking across multiple IBM quantum devices and quantum algorithms validates the framework’s adaptability to different types of superconducting hardware, highlighting its generality and potential for broad adoption in applications. By bridging the gap between theoretical models and experimental reality, our data-efficient approach provides a critical enabling technology for developing the next generation of effective error mitigation techniques and noise-aware compilers, accelerating the realization of practical quantum computing applications.
@@
Future work should balance model complexity and computational cost for scaling to larger quantum systems. More parameters enhance accuracy but increase ML optimization time and data requirements, potentially requiring dedicated characterization circuits to prevent overfitting. A promising route is hierarchical noise modeling that partitions the coupling graph into regional subgraphs with region-specific parameters, capturing spatial inhomogeneity without percomponent overparameterization and preserving scalability. Additionally, extending our framework to model leakage, where qubits excite outside of the computational subspace, will be critical. This requires higher-dimensional simulations and often specialized characterization protocols, substantially increasing the computational burden of the ML loop. Although beyond the scope of this initial study, incorporating leakage is essential for further fidelity gains. The ultimate goal is a full integration into quantum compiler toolchains, enabling adaptive, algorithm-specific compilation that dynamically tailors noise models to both hardware’s current state and circuit structure. This would optimize execution by capturing the unique interplay between algorithms and device-specific error channels. Finally, extending to trapped-ion and photonic platforms will establish universal, cross-platform characterization protocols.
@@ Acknowledgment
The paper was written as part of the project “Quantum-based Energy Grids (QuGrids)”, which is receiving funding from the programme “Profilbildung 2022”, an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. The contribution of M. R. and D. A. K. has been supported by the Dieter Schwarz Foundation through the Fraunhofer Heilbronn Research and Innovation Centers HNFIZ. Parts of the data used for this study have been generated in the course of propject QORA funded by by Ministerium für Wirtschaft, Arbeit und Tourismus Baden-Württemberg in the frame of the Competence Center Quantum Computing Baden-Württemberg. We acknowledge the use of IBM Quantum services for this work and to advanced services provided by the IBM Quantum Researchers Program. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v2

@@ III.3 Random circuits
We further benchmark the method through random circuits on ibmq_kolkata. Random quantum circuits were generated using Qiskit’s random_circuit function with systematically varied parameters: qubit count $n\in\{4,...,10\}$ and depth $d=n+3$. The function constructs circuits by randomly sampling single-qubit gates from SU(2) and placing two-qubit gates between random qubit pairs within the depth constraint, with terminal measurements on all qubits. The results show a substantial mean $D_{H}$ improvement of 43.60% in model fidelity for training and 45.90% for prediction (see Fig. 7(b) in Appendix A). We perform three independent optimization processes and achieve a mean Hellinger distance reduction of $42.9\%\pm 1.5\%$ for the training set and $46.8\%\pm 2.1\%$ for the prediction set, with a maximum reduction of 63% observed for a circuit in the prediction set. This demonstrates that our method’s high performance is a statistically significant result, making it a reliable tool for practical application-aware hardware characterization.
@@
III.4 Application-aware noise characterization
The results show that the proposed framework achieves high predictive accuracy across three qualitatively different algorithmic families (QAOA, VQE, and random circuits). Crucially, rather than seeking a single universal parameter vector $\bm{\theta}$ that simultaneously describes all algorithm classes on a target device, we adopt an application-aware philosophy: we estimate independent parameter sets for each algorithm-hardware context while retaining a fixed model architecture. This design is physically motivated by the context-dependent nature of noise in superconducting processors. Different algorithms can emphasize different effective error mechanisms through their circuit structure (e.g., depth, routing, and parallelism). For instance, QAOA applied to portfolio-optimization problems can require extensive SWAP-network routing to achieve all-to-all connectivity, whereas VQE with hardware-efficient ansätze typically conforms to the native coupling map, resulting in significantly shallower depths and fewer entangling gates. Since our parameterization captures duration-aware dephasing and correlated errors, we do not expect a parameter set calibrated on high-depth QAOA circuits to accurately predict low-depth VQE fidelity, or vice versa. Generality is thus achieved at the framework level: the same workflow applies across devices and algorithms, while the effective parameters adapt to the specific error dynamics of each context.
While cross-algorithm parameter transfer is not a prerequisite for deployment, quantifying it would elucidate the fundamental trade-off between intrinsic device physics and application-specific error manifestations. Investigating the predictive validity of a model trained on one algorithm family (e.g., QAOA) when applied to another (e.g., VQE) helps distinguish between static hardware properties and the effective error dynamics induced by specific circuit designs. Future work utilizing transfer learning could quantify the adaptation cost between these contexts, ultimately determining whether the fidelity benefits of application-aware calibration justify the overhead of independent training or if approximate universality offers a viable compromise for large-scale systems.
@@ IV Discussion and Conclusion
This work develops a novel parameterized noise model with circuit-size-independent complexity. Leveraging machine learning-driven optimization, we repurpose data from a few routine circuit executions to construct data-efficient and prediction accurate noise models for superconducting quantum processors. Crucially, our framework prioritizes application-aware characterization: we estimate effective noise parameters independently for each algorithm family on the target hardware rather than seeking a universal parameter set across all algorithm families. This design captures the distinct interaction between the algorithm and hardware, achieving high predictive accuracy with no additional dedicated calibration circuits or QPU jobs beyond the workload executions themselves. By systematically searching complex, multi-parameter error spaces, our approach consistently produces models that are significantly more predictive than default ones obtained from hardware vendors. We demonstrate a reduction of up to 65% in the Hellinger distance, a metric that quantifies the divergence between predicted and experimental measurement-outcome (bitstring) probability distributions, confirming that our optimized models produce output distributions that more accurately represent the behavior of algorithms executed on real quantum hardware. This significant improvement in predictive accuracy is essential for a reliable simulation of quantum circuits on noisy quantum devices, providing a foundation for more accurate noise-aware quantum circuit compilation by enabling informed optimization of qubit selection, SWAP gate insertion, and other compilation passes.
@@
A key advantage of our approach is its efficiency and practicality. It repurposes existing benchmarking or application-specific execution data, avoiding the need for extensive characterization protocols such as quantum process tomography. This reusability reduces experimental overhead, enabling practitioners to leverage prior executions to improve algorithm testing and error mitigation protocols while iteratively refining noise models as additional data become available. Comprehensive benchmarking across multiple IBM quantum devices and quantum algorithms validates the framework’s adaptability to different types of superconducting hardware, highlighting its generality and potential for broad adoption in applications. By bridging the gap between theoretical models and experimental reality, our data-efficient approach provides a critical enabling technology for developing the next generation of effective error mitigation techniques and noise-aware compilers, accelerating the realization of practical quantum computing applications.
@@
Future work should address scaling challenges through hierarchical noise modeling. Large systems may require richer parameterizations to enhance model accuracy, which can increase ML optimization time and data requirements. Partitioning the device coupling graph into regional subgraphs with region-specific parameters would capture spatial inhomogeneity in larger systems without per-component overparameterization, representing a valuable future direction. Additionally, investigating transfer learning across algorithmic families could quantify the adaptation cost between algorithmic contexts, helping determine if approximate universality is a viable compromise for large-scale systems. Moreover, incorporating leakage errors, where qubits excite outside of the computational subspace, represents a critical extension. This requires higher-dimensional simulations and often specialized characterization protocols, substantially increasing the computational burden of the ML loop. Although beyond the scope of this initial study, incorporating leakage is essential for further fidelity gains.
Our pairwise readout correlation model captures nearest-neighbor effects but neglects long-range correlations; small cluster-wise confusion matrices tied to resonators or readout lines could provide systematic improvements. Benchmarking dense algorithms like QAOA on utility-scale processors would complement our VQE results on 156-qubit hardware, revealing how the framework interacts with large-scale qubit mapping strategies. Additionally, extending to alternative qubit modalities represents essential future work. While our experiments focus on superconducting devices, the methodology is architecture-agnostic. Adaptation requires replacing the physics-informed ansatz $\mathcal{N}(\bm{\theta})$ with platform-specific mechanisms. For trapped-ion processors [bruzewicz2019trapped, moses2023arace], for instance, a tailored ansatz would capture dominant error sources in Mølmer–Sørensen gates, such as motional mode heating and laser intensity fluctuations. The ultimate goal is a full integration into quantum compiler toolchains, enabling adaptive, algorithm-specific compilation that can dynamically tailor noise models to both the hardware’s current state and circuit structure. This would optimize execution by capturing the unique interplay between algorithms and device-specific error channels.
@@ Acknowledgment
The paper was written as part of the project “Quantum-based Energy Grids (QuGrids)”, which is receiving funding from the programme “Profilbildung 2022”, an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. The contribution of M. R. and D. A. K. has been supported by the Fraunhofer Heilbronn Research and Innovation Centers HNFIZ. Parts of the data used for this study have been generated in the course of project QORA funded by by Ministerium für Wirtschaft, Arbeit und Tourismus Baden-Württemberg in the frame of the Competence Center Quantum Computing Baden-Württemberg. We acknowledge the use of IBM Quantum services for this work and to advanced services provided by the IBM Quantum Researchers Program. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.